OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (5, -5) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

Polar coordinates are (r, Θ), where \[r=\sqrt{x^2+y^2}\] and \[\theta=\tan^{-1} \frac{ y }{ x }\]

OpenStudy (anonymous):

can you calculate a value for r?

OpenStudy (anonymous):

\[r = 5\sqrt{2}\]

OpenStudy (anonymous):

yes. What about Θ?

OpenStudy (anonymous):

\[\Theta = \frac{ 3\pi }{ 4 }, \frac{ 7\pi }{ 4 }\]

OpenStudy (anonymous):

sort of. First off the angles are in degrees. (5, -5) is in the 4th quadrant, and 7π/4 or 315° is the angle in the 4th quadrant. So one pair is (5√2, 315°).

OpenStudy (anonymous):

To find the second pair, you can use the 2nd quadrant angle (basically a 180° rotation), but then you have to reflect the r over the origin. So the 2nd pair is (-5√2, 135°)